If the outliers are not included, what is the mean of the data set?
72,82,3,55,62,73,90,80,151,59,46,31
A.59
B.65
C.67
D.73
1
Expert's answer
2014-03-04T12:22:05-0500
The mean (or average) of a set of data values is the sum of all of the data values divided by the number of data values.
Therefrom,
mean=(72+82+3+55+62+73+90+80+151+659+46+31)/12=804/12=67.
The right answer is C.